# GraphToxin: Reconstructing Full Unlearned Graphs from Graph Unlearning

### v1

@@ 5.1 Graph Unlearning
According to Said et al. [surveygraphunlearning], existing graph unlearning methods can be categorized into exact graph unlearning and approximate graph unlearning.
@@ 5.1.1 Exact Graph Unlearning
GraphEraser pioneered the first graph unlearning framework [graph_unlearn]. Specifically, it partitions the original training graph into disjoint shards and then trains a separate GNN on each shard. Upon a new unlearning request, it directly deletes the node from the corresponding shard and subsequently retrains the shard GNN. Wang et al. [inductive_gu] later adapted this framework into an inductive setting.
@@ 5.1.2 Approximate Graph Unlearning
GIF [GIF] utilizes the Graph Influence Function to estimate parameter changes after data removal and then computes a corrective update to these parameters to approximate the state of a model retrained on the reduced dataset. GNNDelete [gnn_bench_nas] optimizes two key objectives: deleted edge consistency, which scrubs the deleted influence from both model weights and neighboring representations, and neighborhood influence, which ensures that retained model knowledge is unaffected after deletion. MEGU [MEGU] introduces a mutual evolution unlearning framework, which simultaneously optimizes a predictive module and an unlearning module to generate effective and reliable unlearning performance. ETR [ETR] proposes a two-stage training-free unlearning framework. In the Erase phase, ETR modifies model parameters to mitigate the impact of deleted samples and their propagated influence on connected nodes. In the Rectify phase, it approximates the gradients of the remaining samples to enhance utility after graph unlearning.
5.2 Graph Reconstruction Attack against GNNs
The related work on graph reconstruction attacks against GNNs can be classified into the following four categories:
@@ 5.2.1 Link Stealing Attack
He et al. [link_steal_gnn] introduced the first link stealing attack and provided a comprehensive attack taxonomy using various background knowledge. LinkTeller [link_teller] exploits node features and labels to propose the first query-based link stealing attack via influence analysis in both transductive and inductive settings. LinkThief [linkthief] combines node similarity, generalized structural features, and shadow graphs to assist link membership inference. Recently, TrendAttack [unlearning_inversion_attacks_graph] emerged as the first link stealing attack against graph unlearning to identify the link membership through confidence trend features. However, its attack assumptions are unrealistic in practice. Furthermore, it solely concentrates on an edge removal, whereas our research targets the hardest unlearning setting—node removal, posing substantially greater challenges and actual risks.
@@ 5.2.2 Attribute Inference Attack against GNNs
Duddu et al. [mia_gra_gnn] proposed the first attribute inference attack against GNNs. They exploited the node embeddings of partial nodes and their associated sensitive features to train a supervised attack classifier and ultimately infer the sensitive attributes of the target nodes. Olatunji et al. [aia_gnn] developed three attribute inference attacks that leverage partial sensitive attributes, public non-sensitive attributes and shadow graphs to conduct multiple queries. However, these attacks solely target partial sensitive attributes, rather than full node feature recovery, which is one of the primary focuses of our work.
@@ 5.2.3 Graph Topology Reconstruction Attack
Duddu et al. [mia_gra_gnn] also designed the first graph topology reconstruction attack to exploit a strong correlation between publicly released embeddings and the ground-truth graph topology. Following the same attack setting, GraphMI [graphmi] and Zhang et al. [inference_gnn] reconstructed graph topology with similar graph statistics to the ground-truth through a graph encoder-decoder paradigm. Olatunji et al. [explain_gra] instead employed node features and their explanations for graph topology recovery. Specifically, they aimed to minimize the reconstruction error on node features and explanations and the loss of node-level predictions. However, as previously mentioned, these attacks can only reconstruct graph topology instead of the full graph components. Furthermore, their attack assumptions are unrealistic, preventing generalization to the data-free black-box setting.
@@ 5.2.4 Gradient Inversion Attack
While the above three attack types have broad application scenarios, gradient inversion attacks (GIAs) against GNNs are specifically tailored for graph federated learning. GRAIN [grain] presents the first GIA against GNNs for partial node feature and graph topology recovery. It employs the rank-deficiency of the GNN layers to reconstruct candidate subgraphs, iteratively filters unreliable ones, and finally combines these candidates with a depth-first search strategy for graph recovery. GLG [gradient_attack_gnn] successfully validates that the subgraphs can be reconstructed through gradient matching in both node classification and graph classification tasks. However, these GIAs cannot be trivially adapted to graph unlearning.
@@ 6 Conclusion
In this study, we pioneer the investigation into the vulnerabilities of reconstructing the full unlearned graphs through graph unlearning. We propose GraphToxin, the first full graph reconstruction attack against graph unlearning. Specifically, we introduce a novel curvature matching module to provide fine-grained signals for full unlearned graph recovery in more realistic attack settings. We further extend GraphToxin to multiple node removals under both white-box and black-box settings. Furthermore, we present a rigorous and comprehensive evaluation framework, demonstrating the necessity of diverse evaluation metrics and worst-case analysis. Our attack results from extensive experiments confirm the effectiveness and flexibility of GraphToxin in both white-box and black-box settings for a single and multiple node removals. Notably, our assessment of existing defense mechanisms reveals their general ineffectiveness against this attack, highlighting the urgent need to develop more robust defensive countermeasures.
Ethics Considerations
@@
LLM Usage Considerations
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.
.1 Experimental Validation
.1.1 Datasets
Cora and PubMed are small-scale and medium-scale citation networks with low average node degrees, where nodes represent scientific/medical publications, edges denote citations, and node features carry author’s personal information. Photo is a dense co-purchase graph, where nodes represent products, edges indicate frequent co-purchases, and features encode product reviews. The dataset statistics are shown in Table IV. We follow the classical data split setting in GNN benchmarking [pitfallsgraphneuralnetwork, gnn_bench_nas] to randomly choose $20$ nodes per class as the training set, $30$ nodes per class as the validation set, and the rest as the test set.
TABLE IV: Detailed Graph Statistics
@@
    GraphToxin: For a single node removal, we select $10\%$ random nodes or the top $10\%$ nodes with the highest degrees from the training graph as attack targets, and report the average results throughout our paper. Please note that attackers do not know the exact location of these nodes. They can only obtain the gradient difference between the original and unlearned GNNs. We set the Fisher coefficient as $1$, the Laplacian coefficient as 1e51e\!-\!5, and use AdamW [adamw] as the optimizer. Since multiple restarts to the optimizer (e.g, L-BFGS) from different starting points can significantly improve the reconstruction quality, for a fair comparison, we strictly follow the strategy in Sinha et.al. [gradient_attack_gnn] for all experiments and only report the attack results with a single run of the optimizer. Similarly, we strictly set the attack learning rate as $0.01$ for all experiments. Notably, better attack results can be achieved by carefully tuning this hyperparameter. The optimization runs for up to $10,000$ iterations, but most attacks stopped earlier. Following Geiping et.al. [first_gradient] and Hu et.al. [first_unlearn_attack], we apply a step size decay after $3/8$, $5/8$, $7/8$ of total iterations, using the multiplicative factor of $0.5$.
.1.3 Original and Unlearning Performance
See Table V.
TABLE V: The Performance of the Original and Unlearned GNNs for A Single Node Removal
@@
.1.4 Attack with GNN Backbones
See Table VI.
TABLE VI: GraphToxin for A Single Node Removal with Different GNN Backbones (GU: Retraining). Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).
@@
.1.5 Impact Factor - Deletion Graph Size
See Figure 5.
Figure 5: Impact of the Full Unlearned Graph Size
.1.6 Impact Factor - Node Homophily
See Figure 6.
Figure 6: Impact of Node Homophily
.1.7 Multiple Node Removal
We examine the efficacy of GraphToxin when removing $5,10,15$ nodes simultaneously from the training graph, representing a more challenging unlearning scenario. For a fair comparison, all hyperparameters are kept identical to the single-node removal experiments. We consider two unlearning settings: in the random case, we randomly choose $5$ disjoint sets for each removal size. While in the worst case, we select the top $25,50,75$ nodes with the highest degrees and then partition them into $5$ non-overlapping pairs based on their degree ranking. The average results are reported in Table VII. Notably, more competitive performance can be achieved by thorough parameter tuning.
Interestingly, the unlearning performance of Retraining in the worst-case multiple removal setting improves in accuracy but slightly decreases in AUC when more nodes are unlearned. As for attack performance, similarly as before, GraphToxin retains its superiority across all metrics in multiple node removals. When the number of deleted nodes increases, the attack performance on the reconstruction-related metrics improves, this is reflected in a reduction in NRMSE (all cases) and embedding distance (random case), and an increase in PMGK (worst case). In contrast, the trends for the performance-level metrics diverge in these two scenarios. For the worst-case removal, the attack’s performance steadily degrades as more nodes are removed. The random case, however, exhibits a non-monotonic trend, where the performance first improves and then declines. Though targeting multiple node removals results in a decrease in overall attack performance compared to the single-node scenario, GraphToxin remains a potent threat.
TABLE VII: GraphToxin for Multiple Node Removals Based on Retraining (Backbone: GCN). Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).
@@
.1.8 Black-box Setting
We also assess the performance of GraphToxin under the most intractable black-box setting. Inspired by Zhuang et al. [data_free_msa_gnn], we utilize a GCN with $2$ layers and $256$ hidden units as a copycat, and a $3$-layer MLP with $128$ and $256$ hidden units as a full-parameter graph generator. We set the query graph size as $250$ with $32$ feature dimensions. During model stealing, each query involves training the generator $2$ times with a learning rate of 1e61e\!-\!6 and training the copycat $5$ times with a learning rate of $0.001$. The total number of queries is set to $100$. The stealing performance of the original GNN is $79.40\%$ in accuracy and $96.12\%$ in AUC while the stealing performance of Retraining for both a single and multiple node removals improves in accuracy and slightly declines in AUC. For GraphToxin, we use the exactly same hyperparameter settings as before. As for the newly added semantic calibration module, we set its coefficient as $50$ for the single node removal and $5000$ for the multiple node removal, as the latter introduces much more noises than the single one. We finally report the average results in Table VIII.
Single Node Removal The black-box GraphToxin demonstrates its superiority when targeting a single node removal. Since we introduce the semantic calibration module, the overall attack performance in the performance-level metrics is competitive or even superior in the random case, while the overall attack performance in the reconstruction-related metrics declines. A plausible explanation is that the black-box GraphToxin enforces the semantic calibration for full unlearned graph recovery, which might push some nodes cross the model’s decision boundary to obtain the correct prediction but ignore the actual distance in the feature or embedding space. However, the degradation in such metrics remains acceptable.
Multiple Node Removal Under the black-box multi-node removal setting, GraphToxin remains superior to all baselines. However, compared to the single-node setting, its overall efficacy in the performance-level metrics declines. In contrast, the attack results of the reconstruction-related metrics are mixed. The performance worsens in PMGK (both cases) and embedding distance (random case) while the attack significantly improves in NRMSE (both cases) and embedding distance (worst case).
TABLE VIII: Black-box GraphToxin for A Single and Multiple Node Removals Based on Retraining. Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).

### v2

@@ 5.1 Graph Unlearning
Existing graph unlearning methods are commonly categorized into exact unlearning and approximate unlearning [surveygraphunlearning].
@@ 5.1.1 Exact Graph Unlearning
GraphEraser pioneered the first graph unlearning framework [graph_unlearn]. This framework partitions the original training graph into disjoint shards and trains a separate GNN on each shard. Upon receiving a new unlearning request, the target node is removed from its associated shard, followed by retraining of the corresponding shard model. Wang et al. [inductive_gu] subsequently extended this framework to inductive settings.
@@ 5.1.2 Approximate Graph Unlearning
GIF [GIF] utilizes graph influence functions to estimate parameter changes induced by node deletion and subsequently applies corrective updates to model parameters to approximate the state of a model retrained on the remaining graph. GNNDelete [gnn_bench_nas] formulates unlearning as an optimization problem with two objectives: deleted edge consistency, which removes the influence of deleted edges from model parameters and neighboring representations, and neighborhood influence preservation, which ensures that retained model knowledge is unaffected after deletion. MEGU [MEGU] introduces a mutual evolution framework that jointly optimizes a predictive module and an unlearning module to generate effective and reliable unlearning performance. ETR [ETR] proposes a training-free two-stage unlearning framework. In the Erase stage, ETR modifies model parameters to mitigate the impact of deleted nodes and their propagated influence on neighboring nodes. while the Rectify stage approximates the gradients of the remaining nodes to enhance utility after graph unlearning.
5.2 Graph Reconstruction Attacks
Existing graph reconstruction attacks against GNNs can be broadly categorized into four groups based on their attack objectives and assumptions.
@@ 5.2.1 Link Stealing Attack
Link stealing attacks aim to infer the existence of edges in a target graph. He et al. [link_steal_gnn] introduced the first link stealing attack and provided a comprehensive attack taxonomy using various background knowledge. LinkTeller [link_teller] exploits node features and labels to propose the first query-based link stealing attack via influence analysis in both transductive and inductive settings. LinkThief [linkthief] combines node similarity, generalized structural features, and shadow graphs to assist link membership inference. Recently, TrendAttack [unlearning_inversion_attacks_graph] emerged as the first link stealing attack against graph unlearning to identify the link membership through confidence trends. This line of work focuses on edge-level removal, however, node removal introduces substantially broader structural changes and poses greater reconstruction challenges.
@@ 5.2.2 Attribute Inference Attack against GNNs
Attribute inference attacks seek to recover sensitive attributes rather than graph structure. Duddu et al. [mia_gra_gnn] proposed the first attribute inference attack against GNNs. They exploited node embeddings of partial nodes and their associated sensitive features to train a supervised attack classifier for inferring sensitive attributes of the target nodes. Olatunji et al. [aia_gnn] developed three attribute inference attacks that leverage partial sensitive attributes, public non-sensitive attributes and shadow graphs to conduct multiple queries. These attacks solely target sensitive attributes, rather than full node feature recovery, which is a central focus of our work.
@@ 5.2.3 Graph Topology Reconstruction Attack
Graph topology reconstruction attacks aim to recover the adjacency structure of the target graph. Duddu et al. [mia_gra_gnn] designed the first graph topology reconstruction attack to exploit a strong correlation between publicly released embeddings and the ground-truth graph topology. Building on the same attack setting, GraphMI [graphmi] and Zhang et al. [inference_gnn] reconstructed graph topology with graph statistics similar to the ground-truth through a graph encoder-decoder architectures. Olatunji et al. [explain_gra] employed node features and their explanations to guide topology recovery by jointly minimizing reconstruction errors on node features and explanations and prediction losses. These methods reconstruct only graph topology and depend on assumptions that require feature or explanation access, thereby restricting their applicability and failing to generalize to fully data-free black-box settings.
@@ 5.2.4 Gradient Inversion Attack
While the above three attack types have broad application scenarios, gradient inversion attacks against GNNs are specifically tailored for graph federated learning. GRAIN [grain] presents the first gradient inversion attack for partial node feature and graph topology recovery. It initially exploits rank deficiencies in GNN layers to reconstruct candidate subgraphs, then iteratively filters unreliable ones, and finally combines these candidates with a depth-first search strategy for graph recovery. GLG [gradient_attack_gnn] demonstrates that gradient matching enables recovery of subgraphs in both node and graph-level tasks. However, these attacks assume access to client-side gradients and cannot be trivially adapted to graph unlearning.
@@ 6 Conclusion
In this study, we pioneer the investigation into the vulnerabilities of reconstructing the full unlearned graphs through graph unlearning. We propose GraphToxin, the first full graph reconstruction attack against graph unlearning. Specifically, we introduce a novel curvature matching module to provide fine-grained signals for full unlearned graph recovery. We further extend GraphToxin to multiple-node removal under both white-box and black-box settings. Furthermore, we present a comprehensive evaluation framework, demonstrating the necessity of worst-case evaluation and multi-level evaluation metrics. Extensive experiments demonstrates the effectiveness and flexibility of GraphToxin across both white-box and black-box settings for single and multiple-node removal. In addition, our assessment of existing defense mechanisms shows that they provide limited protection and may exhibit unintended effects against this attack, highlighting the urgent need to develop more robust defenses for graph unlearning.
APPENDIX
Ethical Considerations
@@
Open Science
In accordance with the CFP’s Open Science guidelines, we release the artifacts required to reproduce our experiments, including the datasets, code, implementation details, and evaluation pipeline. To preserve anonymity during review, these materials are hosted in an anonymous repository https://anonymous.4open.science/r/GraphToxin-A854, and will be migrated to a public GitHub repository upon acceptance. This artifact release is intended to facilitate further investigation into vulnerabilities in graph unlearning and to inspire future work on robust defenses.
Appendix A Theorem Assumptions and Proof
Assumption 1 (Empirical Risk Function).
We assume that $\mathcal{L}(\cdot)$ is twice continuously differentiable and the Hessian $H=\nabla^{2}\mathcal{L}(\theta)$ is invertible and symmetric w.r.t. parameter $\theta$. Following previous studies [hessian, hessian_influence], we add a damping term $\lambda I$ to $H$ if $\mathcal{L}(\cdot)$ is non-convex.
Assumption 2 (Small Parameter Change).
We empirically observe that this change is small, as reflected in our experimental results. This observation is consistent with the objective of graph unlearning and is theoretically bounded in prior work [GIF, chien2022certified].
Under the above assumptions, the first-order Taylor expansion of the gradient $\nabla\mathcal{L}(\cdot)$ around a stationary point $\theta$ is:
$\begin{split}\nabla\mathcal{L}(\theta^{\ast})\approx\nabla\mathcal{L}(\theta)+\nabla^{2}\mathcal{L}(\theta)(\theta^{\ast}-\theta)\end{split}$ (14)
Gradient-based GU-FGRA initializes a dummy graph and iteratively optimizes it to recover a final graph whose induced gradients explain the observed parameter changes [first_unlearn_attack]. Hence:
$\Delta\tilde{\nabla}\approx\theta^{\ast}-\theta$ (15)
where $\Delta\tilde{\nabla}$ denotes the parameter change induced by the dummy graph. Since the original model is already trained and released, its parameters $\theta$ are fixed. Without loss of generality, GU-FGRA aims to recover parameters $\tilde{\theta}$ such that $\Delta\tilde{\nabla}=\tilde{\theta}-\theta$.
Based on Equation (14) and Equation (15), we provide the proof of Theorem 1 below:
Proof. Based on Equation (14), we get $\Delta\nabla\approx H(\theta^{\ast}-\theta)$. Since H is invertible, $\theta^{\ast}-\theta\approx H^{-1}\Delta\nabla$. As for Equation (15), under the same linearized relation in Equation (14), $\tilde{\theta}-\theta\approx H^{-1}\Delta\tilde{\nabla}$. Therefore, the attack goal is transformed to minimize the difference between $H^{-1}\Delta\nabla$ and $H^{-1}\Delta\tilde{\nabla}$. Formally, $H^{-1}(\Delta\tilde{\nabla}-\Delta\nabla)\approx 0$. Take the $H^{1/2}$ to the left-hand and right-hand side, then we get $H^{-1/2}(\Delta\tilde{\nabla}-\Delta\nabla)\approx 0$. Consider the squared L2 norm, the previous approximation equals to $(\Delta\tilde{\nabla}-\Delta\nabla)^{T}H^{-1}(\Delta\tilde{\nabla}-\Delta\nabla)$, hereby completing the proof.
Appendix B Experimental Validation
B.1 Datasets
Cora and PubMed are small-scale and medium-scale citation networks with low average node degrees, where nodes represent scientific/medical publications, edges denote citations, and node features carry author’s personal information. Photo is a dense co-purchase graph, where nodes represent products, edges indicate frequent co-purchases, and features encode product reviews. The dataset statistics are shown in Table 4. We follow the standard data split setting commonly used in GNN benchmarks [amazon_photo, gnn_bench_nas], randomly selecting $20$ nodes per class for training, $30$ nodes per class for validation, and using the remaining nodes for testing.
Table 4: Detailed Graph Statistics
@@ B.2 Implementation
    GraphToxin: For single-node removal, we randomly sample 10% of training nodes or select the top 10% highest-degree nodes as attack targets, and report their average results throughout our paper. The attacker does not have access to the locations of the removed nodes and only observes the gradient difference between the original and the unlearned GNNs. Unless otherwise specified, the Fisher coefficient is set to $1$ and the Laplacian coefficient is set to 1e51e\!-\!5. AdamW [adamw] is used as the optimizer. Since multiple restarts to the optimizer (e.g, L-BFGS) from different starting points can significantly improve the reconstruction quality, for a fair comparison, we strictly follow the configuration in Sinha et.al. [gradient_attack_gnn] for all experiments and only report the attack results with a single run of the optimizer. In addition, we strictly set the attack learning rate as $0.01$ for all experiments. Following Geiping et al. [first_gradient] and Hu et al. [first_unlearn_attack], each optimization is run for up to $10,000$ iterations, with most attacks converging earlier, and a step size decay is applied after $3/8$, $5/8$, $7/8$ of total iterations, using the multiplicative factor of $0.5$.
B.3 Original and Unlearning Performance
Table 5 shows that all graph unlearning methods exhibit comparable performance across the three GNN backbones, with variations within $1\%$. As for GNN backbones, the GCN backbone achieves the highest performance on Cora and PubMed while SGC performs best on Photo, with differences of less than 2% compared to GCN. Across unlearning methods, ETR preserves model utility more effectively on Cora and PubMed, while Retraining exceeds other unlearning methods with GCN and SGC backbones. In contrast, MEGU yields the strongest results with GraphSAGE on Photo. Regarding removal settings, removing the $10\%$ of the highest-degree nodes results in larger benign performance degradation compared to the random-node removal, which is amplified on sparse and smaller-scale graphs.
Table 5: The Performance of the Original and Unlearned GNNs for A Single Node Removal
@@ B.3 Original and Unlearning Performance
B.4 Attack with GNN Backbones
As shown in Table 6, GraphToxin consistently retains its superiority and generalization across all GNN backbones and graph datasets. Taken together, these results suggest that there is no universally most vulnerable GNN due to the complex interactions among model architecture, graph sparsity and homophily. We provide the following plausible explanations: (1) GCN aggregates information over local neighborhoods, which makes node representations strongly dependent on features propagated from their local neighborhoods. This propagation can provide more informative gradient signals for recovering underlying graph structure. However, on dense graphs, rapid neighborhood expansion can lead to over-smoothing, weakening node-specific signals. As a result, GCN tends to achieve more competitive performance on small-scale and sparse graphs such as Cora. (2) GraphSAGE relies on random neighbor sampling rather than full neighborhood aggregation. Consequently, gradients are computed from incomplete and randomly varying local structures, which might obfuscate the gradient-based guidance for recovery, particularly on sparse graphs. While on dense graphs such as Photo, structural redundancy can offset the randomness introduced by neighbor sampling, enabling GraphToxin with the GraphSAGE backbone to maintain strong reconstruction performance. (3) SGC employs linear multi-hop feature propagation. Although its linearity offers a more direct and simpler proxy for full graph recovery, it can suppress fine-grained and node-specific details. This dual effect might explain why SGC performs well in some metrics on PubMed and Photo, yet exhibits degraded performance on Cora.
Table 6: GraphToxin for A Single Node Removal with Different GNN Backbones (GU: Retraining). Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).
@@ B.4 Attack with GNN Backbones
B.5 Impact Factor - Unlearned Graph Size
From Figures 3 and 6, RNMSE remains consistently low with minimal variation, indicating its insensitivity to the unlearned graph size. PMGK generally decreases as the graph size increases. An exception is observed on Photo with the GCN backbone, where PMGK rebounds when the deletion graph becomes sufficiently large. ED and PWD also exhibit a decreasing trend with increasing unlearned graph size. In addition, attack accuracy and fidelity show noticeable oscillations. These variations arise from single-node removal, where each target node induces an unlearned graph of different sizes and structural complexity. The results are then sorted by the size of the unlearned graph. Heterogeneity across target nodes leads to varying reconstruction difficulty, which manifests as fluctuations in the plotted results.
Figure 6: Impact of the Full Unlearned Graph Size and Node Homophily
B.6 Impact Factor - Node Homophily
Consistent with the analysis on unlearned graph size, RNMSE remains largely invariant to changes in the homophily of unlearned nodes. However, other metrics exhibit clearer correlations with node homophily: PMGK as well as attack accuracy and fidelity are positively correlated with homophily, whereas ED and PWD are negatively correlated. Overall, these observations indicate that node homophily may serve as a more informative indicator than unlearned graph size for characterizing the vulnerability of GU-FGRA. We leave the deeper investigation for future work.
B.7 Multiple Node Removal
In this section, we examine the effectiveness of GraphToxin under multiple-node removal, where $5$, $10$, and $15$ nodes are simultaneously removed from the training graph, respectively. For a fair comparison, all hyperparameters are kept identical to those in single-node removal. Under the random removal setting, we sample $5$ disjoint node sets with the same removal size, while in the worst-case removal setting, we select the top $25,50,75$ highest-degree nodes and partition them into five non-overlapping sets based on their degree ranking. The average results are reported in Table 7.
Under the worst-case multiple-node removal, the unlearning performance of Retraining improves in accuracy but slightly decline in AUC when more nodes are removed. In terms of attack performance, GraphToxin consistently outperforms other baselines. As the number of removed nodes increases, reconstruction-related metrics generally yield better results. In contrast, performance-level metrics display a non-monotonic pattern, which initially increase and then slightly degrade. Additionally, the random-case and worst-case scenarios exhibit complementary strengths, with neither consistently dominating the other across all metrics. This observation highlights the necessity of considering both scenarios for a comprehensive evaluation.
Table 7: GraphToxin for Multiple Node Removals Based on Retraining (Backbone: GCN, Dataset: Cora). Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).
@@ B.7 Multiple Node Removal
B.8 Black-box Setting
Setup We now assess the performance of GraphToxin under a data-free, black-box setting. Following the configuration in Zhuang et al. [data_free_msa_gnn], we utilize a GCN with $2$ layers and $256$ hidden units as a surrogate model, and a $3$-layer MLP with $128$ and $256$ hidden units as a full-parameter graph generator. We set the query graph size as $250$ with $32$ feature dimensions. During model stealing, each query encompasses training the generator $2$ times with a learning rate of 1e61e\!-\!6 and training the surrogate $5$ times with a learning rate of $0.001$. The total number of queries is set to $100$. The stealing performance of the original GNN is $79.40\%$ in accuracy and $96.12\%$ in AUC. After unlearning, the stealing performance of the unlearned GNN under both single and multiple-node removal improves in accuracy while slightly declines in AUC. As for GU-FGRA, we adopt the same hyperparameter settings as in the white-box experiments. For the semantic calibration module, we set its coefficient as $50$ for single-node removal and $5000$ for multiple-node removal, as the latter introduces substantially higher noise. Average results are reported in Table 8.
Single Node Removal We observe that black-box GraphToxin consistently outperforms all baselines under single-node removal. Since we introduce the semantic calibration module to enforce the label alignment, the overall attack performance in the performance-level metrics is competitive, although the performance in reconstruction-related metrics degrades relative to the white-box setting as expected. A plausible explanation is that black-box GraphToxin utilizes the semantic calibration during recovery, which might shift some nodes across the model’s decision boundary to obtain the correct predictions but ignore the actual distance in feature or embedding space. Overall, the degradation in such metrics remains acceptable.
Multiple Node Removal Under the black-box multi-node removal setting, GraphToxin still exceeds other baselines. Compared to single-node removal, performance-level metrics decline when more nodes are unlearned. However, the attack results of the reconstruction-related metrics are mixed. PMGK and ED decrease, while NRMSE improves in both random and worst-case node removal scenarios.
Table 8: Black-box GraphToxin for A Single and Multiple Node Removals Based on Retraining (Dataset: Cora). Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).